Two faulty tennis ball machines start to shoot balls from opposite sides of a 25
meter by 10 meter tennis court. The green ball machine shoots green balls that
stop on the court 5 meters to 20 meters from the green machine's side. The
yellow ball machine shoots yellow balls that stop on the court 2 meters to 16
meters from the yellow machine's side. Find the area of the tennis court that has
balls of either color on it.

Found 2 solutions by hudikamo, ikleyn:
Answer by hudikamo(4)

The Yellow Machine shoots balls over a distance of 14 metres (16-2), 9-23 m away from the green machine.
The Green Machine shoots balls over a distance of 15 metres (20-5), 5-20 m away from itself.
The overlapping area is defined from the Green Machine's nine metre line and 20 metre line, 11 metres wide.
The width of the court is 10 m. Total Area = 11*10 = 110 m^2

Answer by ikleyn(52776)

.
Two faulty tennis ball machines start to shoot balls from opposite sides of a 25
meter by 10 meter tennis court. The green ball machine shoots green balls that
stop on the court 5 meters to 20 meters from the green machine's side. The
yellow ball machine shoots yellow balls that stop on the court 2 meters to 16
meters from the yellow machine's side. Find the area of the tennis court that has
balls of either color on it.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


The balls cover the area from min(5, 2) = 2 meters to max(20,16) = 20 meters.

The length of this area is 20-2 = 18 meters.

Regarding the width of the area, the problem keeps silence about it,
but we can accept/admit that the width is equal to the width of the tennis court of 10 meters.

Thus the covered area is  (or can be estimated as)  18 * 10 = 180 square meters.         ANSWER


Solved.